9. You are standing on a sheet of ice that covers the football stadium
parking lot in Buffalo; there is negligible friction between your feet and
the ice. A friend throws you a 0.400-kg ball that is traveling horizontally
at 10.0 m/s. Your mass is 70.0 kg. a) If you catch the ball, with what
speed do you and move afterwards? b) If the ball hits you and bounces
off your chest, so afterwards it is moving horizontally at 8.0 m/s in the
opposite direction, what is your speed after the collision?
1
Expert's answer
2021-10-11T08:54:43-0400

a)


"0.4(10)=(0.4+70)v_2\\v_2=0.0568\frac{m}{s}"

b)


"0.4(10)=(0.4)(-8)+70v_2^B\\v_2^B=0.103\frac{m}{s}"